# Questions: A Taxonomy for Critical Reflection in Machine-Supported Decision-Making

### v1

@@ 4. Taxonomy of Questions for Critical Reflection
Q10 Model Behaviour Would you suggest the same treatment if the patient were 5 years older?
Q1. Questions can address available information, such as data in the form of provided case information. Previous operations on the spine, for example, might be indicated by integers, but this value does not specify how long ago the surgery was performed and at which location of the spine it was carried out - both aspects that can influence the effectiveness of further operations. Questions can help the decision-maker to inspect and contextualise the data provided to give it semantic depth. In addition to ensuring the quality of the data, question can relate to the reliability of the data provided to ensure that inconsistencies, such as measurement errors, are ruled out, or that the data adequately reflect the intended purpose, such as the pain scores self-assessed by patients using a questionnaire.
Q2. The decisive factors for a specific machine recommendation, i.e., relevance of single data points, are another piece of information that can be questioned. XAI methods, such as SHAP (Lundberg and Lee, 2017) or LIME (Ribeiro et al., 2016), make it possible to extract the contribution of the various features to the outcome and rank them accordingly. In order to stimulate critical reflection, it is possible to use that information to generate an evaluative question, like “Is data point $x$ the one to focus on?” or “How relevant is data $x$?”. By questioning the relevance of decisive data points, the decision-maker is encouraged to contextualise the information, e.g., symptom, in a specific case. To do so, the decision-maker may wish to check and validate the machine recommendation by other means, e.g., by conducting additional tests. So while XAI can show which features have contributed to an outcome, domain knowledge is necessary to evaluate whether these features are important. Questioning the relevance of the most contributing data point also raises the question whether there are other factors that are overlooked in the decision. In addition, considering that different XAI techniques can result in different rankings of the contributing effects of the individual features (Saarela and Jauhiainen, 2021), it is important for decision-makers not to take the extracted information provided by XAI at face value. The contradiction between two opposing feature ranks by two different XAI techniques could also lead to a question like “Is $x$ or $z$ more relevant?”. Consequently, questions of this type (Q2) might overlap with Q4 that relates to the causal structure of recommendations.
Q3. Questions can address the concepts of an approach or problematisation, which in the context of joint decision-making translates to the dataset on which the model is built. Cardiovascular diseases, for example, show different symptoms in men and women. Most of the collected data that is used to train DSS, however, stems from male patients, which is why DSS often lead to misdiagnosis in women (Li et al., 2023). The problem in this case is that certain populations or phenomena are overrepresented in data, which leads to a skewed data distribution (Ueda et al., 2024). Besides, data is often only a partial representation of the phenomena under investigation (Broussard, 2019; Binns et al., 2018), as it is impossible to quantify every aspect of the complexity of social life (Birhane, 2021). Data might have been left out at training the model for different reasons, or was not available in the first place. Consequently, questions like “Does the data adequately represent the phenomena under investigation?”, or “Is more evidence necessary?”, can help to validate the functioning and appropriateness of the model and to learn about its limitations. In addition, questions like “Have you considered information x?” could be helpful if certain data is not included in the dataset and is therefore not considered by the model, although it is (now) known that the missing information contributes to the outcome. So compared to Q2, which is more concerned with the question of what information is relevant in an existing dataset, Q3 asks whether the dataset itself contains the relevant information. To create such questions, documentation of the dataset can be very useful (Gebru et al., 2021; Arnold et al., 2019).
Q4. To further assess the interpretation and inferences of the DSS, questions can help to investigate the (apparent) causal structure of provided case information and resulted machine recommendation. Again, based on feature contribution it is possible to ask whether and how the machine recommendation follows from data point $x$, for instance “Does diagnosis $Y$ follow from symptom $x$?”. So although this question type ‘merely’ reformulates a causal explanation, the aim is to help the decision-maker to analyse and evaluate information.
Q5. Questions can relate to the question or problem to which a DSS is trying to provide an answer, i.e., the purpose of the DSS. So similar to Q4 these questions can address the appropriateness of a recommendation. The focus of Q5, however, is on helping the decision-maker to consider alternatives to the recommendation. Different disease have common symptoms, so the physician wants to make sure that another possible diagnosis is not overlooked. For this, a differential diagnosis can be helpful (Mamede and Schmidt, 2023). At the same time, it can be possible to treat the same condition in different ways. Chronic lower back pain, for example, might be treated by surgical or non-surgical interventions (van Hooff et al., 2014), where the feasibility of each option can depend on constraints like available resources or the patient’s preferences (Q7). For example, if the purpose of a DSS is to optimise for full recovery, it might recommend surgery that involves a long stay in hospital; another purpose might be to promote the well-being of the patient, who may prefer to spend as little time as possible in hospital, so that the more suitable treatment could be a combined physical and psychological program. In this way, there is an overlap with question type Q7 that relates to stakeholder preferences. Nevertheless, the focus of Q5 is on whether alternative recommendations should be considered, and on supporting the decision-maker in developing alternative hypotheses and courses of action.
Q6. Questions can address assumptions and expectations of the decision-maker. On the one hand, these questions can help to elicit tacit knowledge and taken for granted assumptions, which might need to be re-evaluated on a case to case basis. This could prevent the decision-maker from jumping to conclusions. On the other hand, questions that address the expectations of the decision-maker can increase their confidence and allow them to arrive at their own decision. In other words, these questions can help to increase the decision-autonomy of the operator. Example questions could be, “Does the recommendation match your assumptions? If so, why (not)?”. General questions allow to take into account domain expertise that can address factors which might not be available in quantified data (Q3). As such, questions could allow the decision-maker to obtain a more holistic picture of the problem that goes beyond static rules implemented in DSS (van Baalen and Boon, 2015) (Q5).
Q7. Questions can relate to different perspectives or points of view. Especially in decision context in which decision-maker and decision-subject are different persons, such as physician and patient, it is important to take into account stakeholder preferences. As such, questions like “Have you considered the preferences of the patient?”, or “Are there certain situational circumstances that prevent the patient from recovering from surgery for several weeks?”, can help to ensure the decision-maker considers other parties involved and treat the patient in a more personalised way. Ideally, enquiring about the patient’s preferences is part of the physician’s routine practice during consultation, yet these questions can function as a form of safeguard.
Q8. Questions can help to anticipate (unintended) consequences of a decision. These questions can elicit forward-looking responsibility of the decision-maker, for example, by asking “Have you considered the implications of administering drug $x$?” While the decision-maker has ideally weighed up the implications, these questions can function as a reminder. Furthermore, with insights from the dataset (Q3) attention could be drawn to certain limitations of a DSS. Compared to Q3, the focus of Q8 is on assessing the impact of a recommendation that may be based on underrepresented data, and supporting the decision-maker in considering appropriate mitigation strategies.
Q9. Questions can take the form of what-to-be that outcome in order to further investigate the space of different alternatives. By creating hypothetical scenarios, the feasibility of desired outcomes can be investigated. So for example, ‘Is it possible to lower level $z$, this could probably make treatment $y$ more effective?” In doing so, the decision-maker is encouraged to consider prior and smaller interventions that could increase the expected effectiveness of a desired treatment. To create such question, feature perturbation and counterfactual explanations are helpful (Karimi et al., 2021; Poyiadzi et al., 2020). In contrast to counterfactual explanations, which provide information about what the outcome would have been if the data had been like this, these questions are about whether it is possible to change certain data points in order to make a particular outcome more likely. For example, a counterfactual explanation would be “If finding $x$ were not present, the condition would be $y$”, whereas the question turns into “Is it possible to lower $x$, so that the likelihood of condition $y$ increases?”. Although these questions ideally lead to more actionable insights, they might imply unfeasible or unattainable changes and actions, e.g., changing the age of a patient. These hypothetical scenarios can nevertheless provide insights into inbuilt associations and thresholds of the DSS decision-boundaries.
Q10. Questions can address assumptions, as mentions before. In addition to the assumptions of the decision-maker, questions can refer to built-in assumptions or rules of model behaviour. For example, if a DSS makes a decision based on age, with a threshold of 50 years, the decision-maker may want to re-evaluate the decision, if the patient is close to that threshold, e.g., 48 years old. An example question could be “Would you recommend the same treatment if the patient were 3 years older, which might decrease its effectiveness?”. These questions might also relate to the relevance of data points (Q1), as single data points often function as proxies for other characteristics. Age, for example, might be an indication for how well the patient will recover from surgery, yet physical fitness might also play a role. This means, focusing only on single data points could be misleading. As such, these questions can also give the decision-maker a better understanding of the causal structure built into the model (Q4). Documentation on model behaviour can be a source from which questions can be derived (Mitchell et al., 2019).
5. Discussion, Future Work, and Concluding Remarks
We have proposed a taxonomy of question types to support the systematic creation of questions that can stimulate critical reflection during machine-supported decision-making. In doing so, we have focused on the medical domain by providing some example questions. We maintain that our taxonomy is transferable to other decision-making domains. Whilst the wording of the questions may change when applied to other contexts, the general structure and elements of critical reflection remain the same. Future work is required to test and validate the generalisability of our taxonomy. Next to that some considerations remain.
The tone of questions is important. For natural language explanations, it was found that the tone and assertiveness has an effect on decision-making (Calisto et al., 2023). Although questions are in general less assertive compared to explanations, it could be assumed that the tone of questions has a similar effect on how or whether the decision-maker engages with the questions. A question could be framed in a more imposing way, such as, “Can you check for symptom $x$?”, rather than in a suggestive way, e.g., “Have you checked for other symptoms?”. To examine the right tone, we have presented Bloom’s taxonomy. We argue that questions should follow the higher levels of critical thinking, that is evaluating, analysing, and creating. As such, it is likely to assume that the questions should be phrased in an open-ended manner. Simple yes-or-no questions could be overlooked or act as checkboxes without eliciting cognitive engagement. Besides, it is not necessarily the case that definite answers have to be found to questions. Rather, the posed question can lead to further questions and thus inquiry of the decision-maker (Paul, 2019). In order to support reflection effectively, it is therefore necessary to create questions in a controlled and systematic way. Consequently, it might be advisable to (currently) refrain from using large language models (LLMs), as the output risks being irrelevant and unreliable (Weidinger et al., 2022). Besides, conversational presentation of information in the form of chatbots might lead to further overreliance (Anderl et al., 2024), as humanised or anthropomorphised chatbots can lead to an illusion of reciprocity and care (Maeda and Quan-Haase, 2024). Future work could nevertheless investigate how to stimulate continuous reflection by processing the answers of the operator and asking follow-up questions. Ideally, an interactive interface would allow the decision-maker to modify data points accordingly. At the same time, however, decision-makers should not be burdened with too many or too frequent questions. In the aforementioned study on cognitive interventions, the cognitive load of these interventions on the decision-maker was reported, which is why participants favoured them less (Buçinca et al., 2021). The right balance of attention and engagement must be found. The operator should be enabled to constructively take the posed question into account, which also depends at least in part on the competences and confidence of the decision-maker. Ultimately, the cognitive burden is unavoidable, and inherent to critical reflection. As John Dewey (1933) mentions on the difficulty of reflective thinking:
yet thinking need not be reflective. For the person may not be sufficiently critical about the ideas that occur to him. He may jump at a conclusion without weighing the grounds on which it rests; he may forego or unduly shorten the act of hunting, inquiring; he may take the first ‘answer’ or solution, that comes to him because of mental sloth, torpor, impatience to get something settled. One can think reflectively only when one is willing to endure suspense and to undergo the trouble of searching (p17).
With this in mind, and given that questions during the decision-making process are likely to reduce efficiency by interrupting the workflow and taking up more time, the design and implementation of a potential ‘reflection machine’ that supports reflection seems less desirable. However, in domains with high-impact decisions, such as healthcare, it should be justified to make a deliberate and thus perhaps more time-consuming decision, as the outcome can have life-changing effects. Considering that DSS recycle past assumptions with potentially harmful consequences (Burrell, 2024), critical reflection becomes increasingly important to promote responsible decision-making. In particular, forward-looking responsibility, which can be strengthened through reflection, ideally leads to a more open and just future.
In this paper we provided a taxonomy of questions aimed at promoting critical reflection on machine recommendations. Careful selection and presentation of these question types in the right circumstances (e.g., when time for reflection is available), and appropriate combination with recommendations and explanations can support the effective human oversight required by the European AI Act.

### v2

@@ A.3. Our Taxonomy applied to Healthcare
Q10 Model Behaviour Would you suggest the same treatment if the patient were 5 years older? Did you consider that the model makes errors in 1 out of 10 cases?
A.4. Our Taxonomy applied to Education
Table A4. Based on our question taxonomy in Table 1, we provide some sample questions that relate to an educational use case where teachers use learning analytics dashboards to plan lessons, track student progress, set learning goals, and create exercises. The following questions were created with the help of three researchers in education and technology, one of whom is a teacher. Relevant questions will vary depending on use case, these questions should thus be taken as inspiration.
ID Element for Reflection Sample Questions
Q1 Case Information Does the data adequately represent the student’s progress? Are there other factors outside the LAD that affect the student’s progress?
Q2 Relevance of Data Does the failed exercise represent the student’s overall performance? Is the recommended exercise relevant to the learning goal? How important is the result of the failed / passed exercise for the overall progress?
Q3 Dataset What (student) data was the model trained on? Which metrics are included in the dataset?
Q4 Causal Structure of Recommendation Does the recommended exercise contribute to the learning goal? How effective is the recommended exercise for the learning goal?
Q5 Alternatives to Recommendation Are there other suitable exercises that the students can do without their devices (what is pedagogically needed for my class)? What other exercise can be done to achieve the learning goal? Should the student work on a learning goal based on their own progress or based on the level of the whole class?
Q6 Assumptions and Expectations of Decision-Maker What are my assumptions about the student? Are my assumptions about the student adequate?
Q7 Stakeholder Preferences Does the student have any special needs? Can the student benefit from other exercises / help? Are there any (extraordinary) circumstances of the student that should be considered?
Q8 Consequences of Decision Does the recommended exercise contribute to the learning goal? What does it do to the emotional well-being of my student if they have to repeat exercises multiple times based on LAD recommendations? What are the downsides of this exercise? What other learning objectives are being used in this exercise that the student could be struggling with?
Q9 Change Intervention Is it possible to increase the student’s motivation to make the learning goal more likely?
Q10 Model Behaviour Does the LAD contain learning theory / methodology X / adhere to our school-wide policy / goal? Is the model’s assumption that the student will not achieve the learning goal if fail two exercises appropriate?